OpenStudy (anonymous):

Help Please!!!! Will give medal!! At the outer edge of a rotating space habitat, 130 m from its center, the rotational acceleration is g. What is the rotational acceleration at a distance of 65 m from the center of the habitat?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@esshotwired

OpenStudy (anonymous):

@ashore0482

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (anonymous):

In a rotating system, the angular velocity is constant among the disc. What we know is the angular acceleration at a distance b from the center and we want to know the angular acceleration at a distance a from the center.\[\alpha = \omega ^{2}r\]Solving for the angular velocity, we get: \[\omega_{outer} = \sqrt{\alpha_{outer}/b}\]This is also true for the inner radius.\[\omega_{inner} = \sqrt{\alpha_{inner}/a}\]However,\[\omega_{inner} = \omega_{outer}\]Therefore,\[\sqrt{\alpha_{inner}/a} = \sqrt{\alpha_{outer}/b}\]Rearranging the equation, we find\[\alpha_{inner}=\alpha_{outer}*(a/b)\]From here it is simply plug-and-chug with: \[\alpha_{outer}=g, a=65m, b=130m\]I will leave the rest to you.

OpenStudy (anonymous):

I understand that the answer is 130m, right? But I really don't understand how you go to it. Is there a simpler formula that could be used?

OpenStudy (anonymous):

From the problem, you already know that the outer distance is 130m and you are solving for rotational acceleration (which is in units of 1/s^2). The problem is utilizing your knowledge that rotational velocity is constant on a spinning disc and that the formula for rotational acceleration (which I used a poorly chosen symbol) is:\[a_{r} = \omega^{2}r\]In my previous response, I used these two concepts to derive a formula for the rotational acceleration at one part of the disc knowing it at another part and the distance from the center at the two points in question.

OpenStudy (anonymous):

*I meant the units are m/s^2, this is rotational, not radial.